# Revisiting Shooting Point Monte Carlo Methods for Transition Path Sampling

### v1

@@ IV Conclusion
We propose a theoretical framework to describe transition path sampling in an extended ensemble including paths and respective shooting indices. Revisiting algorithms that evolve the shooting point from trial to trial, we derive detailed balance equations based on the extended ensemble leading to corrections for aimless and spring shooting in a flexible-length setting. Although the deviation of path ensembles sampled via aimless shooting or spring shooting from the correct path ensemble is system-dependent, a significant overlap in the distributions was found to remain in our experiments. This indicates that previous works may be critically assessed and qualitative conclusions likely still hold. With recent advances in interface-based sampling schemes such as stone skipping Riccardi et al. (2017) and wire fencing Zhang et al. (2023) for transition interface sampling and virtual interface exchange Brotzakis and Bolhuis (2019) for transition path sampling, we hope to see further development of shooting schemes for TPS that require minimal information about the system of interest.
@@ Acknowledgements.
(Dated: August 6, 2024)
@@
So, by reweighting each sampled path with a factor equal to the inverse of the length of the path, the bias in the path ensemble can be corrected. This is only possible when the two-step shifting of the shooting point is followed rigorously, and therefore, no reweighting factor can be derived for the original spring shooting as proposed in SBrotzakis and Bolhuis (2016). Still, when we use the corrected spring method described in the next section, and use the $1/L(X)$ reweighting factor, we do obtain the correct ensemble, as shown in the Main Text.

### v2

@@ IV Conclusion
We have proposed a theoretical framework to describe transition path sampling in an extended ensemble that includes paths and respective shooting indices. Revisiting algorithms that evolve the shooting point from trial to trial, we derived detailed balance equations for the extended ensemble, which led to corrections for aimless and spring shooting in flexible-length settings. While the deviation of path ensembles sampled via aimless shooting or spring shooting from the correct path ensemble is system-dependent, our experiments demonstrate a significant overlap in the distributions, suggesting that previous studies may be critically assessed and qualitative conclusions may remain largely valid. With recent advances in interface-based sampling schemes such as stone skipping Riccardi et al. (2017) and wire fencing Zhang et al. (2023) for transition interface sampling and virtual interface exchange Brotzakis and Bolhuis (2019) for transition path sampling, we anticipate further development of shooting schemes for TPS that require minimal information about the system of interest.
Data Availability
All data that support the findings of this study are available at 10.5281/zenodo.14753554.
@@ Acknowledgements.
Sebastian Falkner1,2, Alessandro Coretti1, Baron Peters3,4, Peter G. Bolhuis5, and Christoph Dellago1,∗
1Faculty of Physics, University of Vienna, 1090 Vienna, Austria.
2Institute of Physics, University of Augsburg, Universitätsstraße 1, 86159 Augsburg, Germany.
3Chemical and Biomolecular Engineering, University of Illinois at Urbana-Champaign, Urbana, Illinois 61801, USA.
4Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, Illinois 61801, USA.
5Van ’t Hoff Institute for Molecular Sciences, University of Amsterdam, PO Box 94157, 1090GD Amsterdam, The Netherlands.
Electronic address: christoph.dellago@univie.ac.at
(Dated: January 28, 2025)
@@
So, by reweighting each sampled path with a factor equal to the inverse of the length of the path, the bias in the path ensemble can be corrected. This is only possible when the two-step shifting of the shooting point is followed rigorously, and therefore, no reweighting factor can be derived for the original spring shooting as proposed in Brotzakis and Bolhuis (2016). Still, when we use the corrected spring method described in the next section, and use the $1/L(X)$ reweighting factor, we do obtain the correct ensemble, as shown in the Main Text.